OpenStudy (anonymous):

3cos^2 x-2 cos x-2=0

OpenStudy (anonymous):

if you let \(y=\cos x\) then \[ 3y^2-2y-2=0 \]You can solve for \(y\), and it should give two solutions. Suppose they are \(y_1\) and \(y_2\). Then you would use \( y_1 = \cos(x)\) and \( y_2 = \cos(x) \) to get solutions for \(x\).